# Theory of the Spin Seebeck effect induced by crystal-field excitations in Tb3Fe5O12

### v2

@@ I Introduction
The spin Seebeck effect (SSE) is a phenomenon of thermoelectric generation created using a bilayer of a metal and a ferromagnet or ferrimagnet [1, 2, 3, 4, 5, 6, 7, 8]. In the longitudinal spin Seebeck effect, a temperature gradient is applied perpendicular to the interface between a metal and a ferromagnet. This induces a spin current in the metal, which is then converted to a charge current by the inverse spin Hall effect [2, 3]. The charge current is therefore observed as a voltage drop in the direction perpendicular to the temperature gradient. The voltage is estimated by calculating the spin current injected in a metal [4, 5, 6, 7] and this spin current is determined by the thermally populated magnons and their polarizations.
@@
Crystal field excitations (CFE) are key to understand the SSE in TbIG. In the past, optical absorption was used to estimate the crystal-field parameters [17]. However, those parameters do not well reproduce the low energy CFEs observed by neutrons, e.g., 3, 6, and 10 meV [13, 14]. This is perhaps not surprising as the lines that can be observed optically are determined by electric dipole transitions whereas neutron scattering employs the magnetic dipole transitions within the lowest multiplet of total angular momentum and thus will have different selection rules for intensity. In addition, in Ref. 17 effects of the exchange with Fe moments were not considered explicitly. Consequently, the crystal-field parameters must be re-examined in light of the new neutron data. It is noteworthy that the crystal-field levels of RE garnets with non-magnetic transition metals, such as terbium aluminum garnet (TAG) and terbium gallium garnet (TGG), with weak interactions between the terbium ions, are essentially the same at each of the twelve Tb sites in a primitive unit cell. In contrast, in TbIG we have to consider the local contribution of each Tb separately due to the magnetic Fe ions. Exchange with iron atoms perturbs the crystal-field levels in a way that may vary from site to site [22, 19, 20, 21, 18]. Therefore, the local coordination of each Tb must be considered to examine the CFE in TbIG.
@@ II Model and Hamiltonian
In the primitive cell of a physical REIG in the cubic phase, the 32 magnetic ions are distributed as 12 RE+3 and 20 Fe3+, i.e., in a ratio of 3 to 5. Of the 20 Fe3+ sites, 12 (three out of five) are on tetrahedral $d$-sites and 8 (two out of five) on octahedral $a$-sites. Tomasello et al. introduced a simple model respecting stoichiometry with a total of eight sites to understand the essence of the magnetic umbrella structure and dynamics within a spin-wave approximation [18]. For the Tb moments, the single-ion anisotropy was modelled as $S_{\alpha}^{2}$, where the mutually orthogonal easy-axes ($\alpha=X,Y,Z$) were deduced from the crystal-field Hamiltonians measured for TGG and TbIG [21, 17]. The competition between these anisotropies and the Tb-Fe exchanges explained the origin of the single magnetic umbrella in three dimensions. The same authors also suggested that, owing to the local RE crystal-fields and the [111] molecular field from exchange with the Fe sites, the $c$-sites could be distinguished in two groups depending on the breaking of the axial symmetry around the easy ($S_{\alpha}$) and hard ($S_{\beta\neq\alpha}$) axes, thus justifying the double umbrella magnetic structures observed in the full cubic TbIG [27]. Furthermore, the model with two-dimensional “flat-umbrellas” of Ref. 28 allowed for an even simpler algebra to further elucidate the general effects of crystal-field induced non-collinearity. Mechanisms such as hybridization between localised and dispersive modes, and chirality switching in conjunction with level repulsion, were emphasized for potential spin current generation in REIGs  [18, 28]. Only very recently, however, it was reported in Ref. 16 a strongly enhanced spin Seebeck effect in TbIG, with Landau-Lifshitz-Gilbert (LLG) spin-dynamics incorporating two groups of inequivalent single-ion anisotropies in the $c$-sublattices [27]. Here we will consider the full local crystal-field potential around each site, but to keep the model tractable, we consider only a single rare-earth site at a time and include the molecular field of the iron atoms. We must introduce a constraint on the iron moments so that they remain in the [111] direction—this is a slight simplification—in practice even the simple model allows for some of the iron atoms to cant with the rare-earths. The iron canting angles are much less than for the rare-earths, because of cancellation from different rare-earth-iron magnetic exchanges.
@@
where the expansion in terms of the Stevens operators ${O}_{q}^{k}({J}_{z},{J}_{\pm})$ account for the $D_{2}$ point-group symmetry of the $c$-sites of Tb garnet crystals [19, 20, 21]. The operators ${J}_{z},{J}_{\pm}$ and thus $H_{\mathrm{CF}}$ in Eq. (2) are defined in a “local coordinate system” convenient for the $D_{2}$ symmetry due to the oxygens surrounding the RE sites (more is discussed in Sec. III below, and further details can be found in Appendix A). Our results here make use of the parameters $\tilde{B}_{q}^{k}$ estimated via neutron spectroscopy on TGG by Wawryńczk et al. [21]. These TGG parameters are listed in the left column of Table 1 together with the corresponding values for TbIG obtained by Hau et al. via infra-red spectroscopy [17].
@@
In the mean field calculation below, the TGG parameters for $H_{\mathrm{CF}}$ in Eq. (2) are used to set the unperturbed CF Hamiltonian for Tb ions in TbIG. Each Tb site is then independently and separately examined using the Hamiltonian in Eq. (1) and one of the the six different coordinate systems exploiting Eq. (2) (details in Appendix A). We underline that our theoretical results for the SSE by CFE in TbIG are obtained by using the TGG parameters, from the left column of Table 1.
TGG, Ref. 21 TbIG, Ref. 17
@@
Table 1: Parameters (in meV) for the Stevens’ CF Hamiltonian in Eq. ((2)). All values are from the experimental literature: neutron spectroscopy on Tb3Ga5O12 (TGG); optical spectroscopy on Tb3Fe5O12 (TbIG) [17, 21]. For TbIG, the $\tilde{B}_{q}^{k}$ Stevens parameters (here reported in meV) are converted from the Wybourne parameters in Ref. 17 (where they are expressed in cm-1).
@@ III Mean field solution of each Tb site
where $n_{\alpha}$ and $z_{\alpha\beta}$ denote, respectively, the site number in a unit cell of the model (see Fig. 1) and the number of neighbouring sites, with $\alpha,\beta=a,d,c$ for the site species. Here, the $i$ index e.g., Eq. (6) indicates one of the six different sublattices of Tb sites, with $i=1,2,\dots,6$ following the notation adopted in Refs. 19, 20. Among these, the sites 1, 2 and 3 (4, 5, and 6) are assigned to the symmetry and g-tensors of the $\mathbf{C}$-sites ($\mathbf{C^{\prime}}$-sites) [23, 24, 26]. The three $\mathbf{C}$-sites ($\mathbf{C^{\prime}}$-sites) are related by 120 degrees rotation around the [111] axis of a garnet crystal.
@@
$\cdot$ Local RE coordinates, $\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{z}_{i}$ (see Appendix A);
@@
These coordinate axes are assigned to represent features of the real garnet crystals, with the Fe moments aligning ferrimagnetically along the [111] of the “laboratory coordinate”—i.e., the azimuth $\mathbf{z}=[111]$ of the “magnetic coordinate”—and the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ symmetries dictating the “local coordinates” for the RE sublattices (see Appendix A[23, 24, 26, 19, 20, 21, 18]. In $H_{\mathrm{MFi}}$, the “magnetic coordinate” system is used even for the Tb moments, for a unified description with the Fe moments. Using Eq. (6), the CF Hamiltonian is rotated from the $i$-th “local coordinate” of the Tb sites into the “magnetic coordinate” convenient for the mean-field model of Fig.1. These rotations are realised by the Wigner matrices ${\cal D}(0,\theta,\phi),{\cal D}(\alpha_{i},\beta_{i},\gamma_{i})$, with $(0,\theta=\arctan\!\!\sqrt{2},\phi=-\pi/4)$ and $(\alpha_{i},\beta_{i},\gamma_{i})$ the Euler angles of the “magnetic” and “local” coordinates, respectively, as defined in the “laboratory coordinates” [19, 20]. The explicit parametrisation of the $\alpha_{i},\beta_{i},\gamma_{i}$ Euler angles and the respective local coordinates $\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{z}_{i}$ for each $i$-th Tb sublattice are given in Appendix A.
@@
In Fig. 2, the summations of six cases for each ion, i.e., $M_{c}^{z}=\sum_{i=1~{}6}M_{c}^{z}(i)$, $M_{d}^{z}=\sum_{i=1~{}6}M_{d}^{z}(i)$, and $M_{a}^{z}=\sum_{i=1~{}6}(2/3)M_{a}^{z}(i)$, are plotted. This corresponds to a half of a primitive cell containing six Tb sites, six $d$ sites and four $a$ sites. Here, the $x,y,z$ indexes refer to the axes of the “magnetic” coordinate system as defined above. Thus, the spins ${\vec{S}}_{a}$ and ${\vec{S}}_{d}$, as calculated in Eq. (1), and the $\vec{J}_{c}$ are expressed in this same coordinate system, as anticipated. Since the $a$ and $d$ moments of the Fe sublattices are assumed to be antiparallel along the azimuth $z$—remember this coincides with the [111] of the “laboratory coordinate system”—these will only have a net $z$ component, i.e., $M_{a}^{x}=M_{a}^{y}=M_{d}^{x}=M_{d}^{y}=0$, whereas all components of the $c$ magnetic moments (Tb sites) can be non-zero below the Néel temperature. After taking the summation, the $x$ and $y$ components of Tb sites are zero, i.e., $M_{c}^{x}=\sum_{i=1~{}6}M_{c}^{x}(i)=0$ and $M_{c}^{y}=\sum_{i=1~{}6}M_{c}^{y}(i)=0$, since three sites among $\mathbf{C}$-site ($\mathbf{C^{\prime}}$-sites) are related by the 120 rotation around the azimuth $z$. We note that this is a simplification of the model: in the real lattice the overall magnetisation along the [111] of the crystal is stable because of the net cancellation of the transverse contribution of the c-sites, and we would also expect induced transverse components on at least some of the Fe sites, as seen in the simplest model [18].
@@
The temperature dependence of the crystal-field levels from the mean field solution of Eq. (3) are plotted in Figs. 4. The different curves can be grouped into the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ results, since the molecular field along the crystallographic [111] has the same projections on the $i=1,2$ and $3$ RE sites and, separately, on the $i=4,5$ and $6$, as discussed in Ref. 18. At the Néel temperature, the molecular field goes to zero and thus the TGG energies observed by Wawryńczk et al. are reproduced [21]. It is noteworthy that, as the molecular field increases for lower temperatures, low energy excitations appear at around 3 meV and 6 meV, as observed by Kawamoto et al. [13].
@@
Figure 5: Magnetic field (B) dependence of the lowest CFE of a $\mathbf{C}$-site at $T=$ 10 K. The horizontal axis $B$ is the applied field in the [111] direction. The inset is the same but for the second CFE. For the $\mathbf{C}$-sites the lowest energy levels are higher in energy and increase with field, (see Appendix B ).
Remarkably, the lowest CFE decreases with increasing the strength of the applied magnetic field. This is in sharp contrast to the magnetic excitation gap of a ferrimagnetic Heisenberg model [10] and illustrates the importance of including the effects of the strong anisotropies for compounds where rare-earth ions have non-zero orbital angular momentum. On the other hand, the second CFE of the $\mathbf{C}$-site (see inset of Fig. 5) increases with the magnetic field. Although the results in Fig. 5 might seem to be at odds with our intuition, the magnetization of the Fe sites resolves this apparent contradiction. The magnetization of the iron sites generates a molecular field with magnitude around 14.4 T. When an external magnetic field is applied in addition to this, the lowest CFE decreases for up to around 15 T of the external magnetic field. This can be verified already at a single-ion level: for the $\mathbf{C}$-sites the CF Hamiltonian from TGG with a magnetic field applied along the crystallographic [111] reproduces this behaviour, whilst for the same conditions for the $\mathbf{C^{\prime}}$-sites the levels all increase with magnetic field strength, closer to intuition based on isotropic models (details can be found in Appendix B). Therefore, one can say that Fig. 5 is reasonable in TbIG.
Reflecting upon these findings, a question arises: Does the low energy CFE contribute to generate a spin current? In the next section, we propose a simple model to examine this question and provide an answer.

### v3

# Theory of the spin Seebeck effect influenced by crystal-field excitations in Tb3Fe5O12

@@ I Introduction
The spin Seebeck effect (SSE) is a phenomenon of thermoelectric generation created using a bilayer of a metal and a magnet, that can be a ferromagnet or ferrimagnet [1, 2, 3, 4, 5, 6, 7, 8] or even an antiferromagnet  [9, 10, 11]. In the longitudinal spin Seebeck effect, a temperature gradient is applied perpendicular to the interface between a metal and a ferromagnet. This induces a spin current in the metal, which is then converted to a charge current by the inverse spin Hall effect [2, 3]. The charge current is therefore observed as a voltage drop in the direction perpendicular to the temperature gradient. The voltage is estimated by calculating the spin current injected in a metal [4, 5, 6, 7, 11] and this spin current is determined by the thermally populated magnons and their polarizations.
@@
Crystal field excitations (CFEs) are key to understand the SSE in TbIG. In the past, optical absorption was used to estimate the crystal-field parameters [20]. However, those parameters do not well reproduce the low energy CFEs observed by neutrons, e.g., 3, 6, and 10 meV [16, 17]. This is perhaps not surprising as the lines that can be observed optically are determined by electric dipole transitions whereas neutron scattering employs the magnetic dipole transitions within the lowest multiplet of total angular momentum and thus will have different selection rules for intensity. In addition, in Ref. 20 effects of the exchange with Fe moments were not considered explicitly. Consequently, the crystal-field parameters must be re-examined in light of the new neutron data. It is noteworthy that the crystal-field levels of RE garnets with non-magnetic transition metals, such as terbium aluminum garnet (TAG) and terbium gallium garnet (TGG), with weak interactions between the terbium ions, are essentially the same at each of the twelve Tb sites in a primitive unit cell. In contrast, in TbIG we have to consider the local contribution of each Tb separately due to the magnetic Fe ions. Exchange with iron atoms perturbs the crystal-field levels in a way that may vary from site to site [25, 22, 23, 24, 21]. Therefore, the local coordination of each Tb must be considered to examine the CFE in TbIG.
@@ II Model and Hamiltonian
In the primitive cell of a physical REIG in the cubic phase, the 32 magnetic ions are distributed as 12 RE+3 and 20 Fe3+, i.e., in a ratio of 3 to 5. Of the 20 Fe3+ sites, 12 (three out of five) are on tetrahedral $d$-sites and 8 (two out of five) on octahedral $a$-sites. Tomasello et al. introduced a simple model respecting stoichiometry with a total of eight sites to understand the essence of the magnetic umbrella structure and dynamics within a spin-wave approximation [21]. For the Tb moments, the single-ion anisotropy was modelled as $S_{\alpha}^{2}$, where the mutually orthogonal easy-axes ($\alpha=X,Y,Z$) were deduced from the crystal-field Hamiltonians measured for TGG and TbIG [24, 20]. The competition between these anisotropies and the Tb-Fe exchanges explained the origin of the single magnetic umbrella in three dimensions. The same authors also suggested that, owing to the local RE crystal-fields and the [111] molecular field from exchange with the Fe sites, the $c$-sites could be distinguished in two groups depending on the breaking of the axial symmetry around the easy ($S_{\alpha}$) and hard ($S_{\beta\neq\alpha}$) axes, thus justifying the double umbrella magnetic structures observed in the full cubic TbIG [30]. Furthermore, the model with two-dimensional “flat-umbrellas” of Ref. 31 allowed for an even simpler algebra to further elucidate the general effects of crystal-field induced non-collinearity. Mechanisms such as hybridization between localised and dispersive modes, and chirality switching in conjunction with level repulsion, were emphasized for potential spin current generation in REIGs  [21, 31]. Only very recently, however, it was reported in Ref. 19 a strongly enhanced spin Seebeck effect in TbIG, with Landau-Lifshitz-Gilbert (LLG) spin-dynamics incorporating two groups of inequivalent single-ion anisotropies in the $c$-sublattices [30]. Here we will consider the full local crystal-field potential around each site, but to keep the model tractable, we consider only a single rare-earth site at a time and include the molecular field of the iron atoms. We must introduce a constraint on the iron moments so that they remain in the [111] direction—this is a slight simplification—in practice even the simple model allows for some of the iron atoms to cant with the rare-earths. The iron canting angles are much less than for the rare-earths, because of cancellation from different rare-earth-iron magnetic exchanges.
@@
where the expansion in terms of the Stevens operators ${O}_{q}^{k}({J}_{z},{J}_{\pm})$ account for the $D_{2}$ point-group symmetry of the $c$-sites of Tb garnet crystals [22, 23, 24]. The operators ${J}_{z},{J}_{\pm}$ and thus $H_{\mathrm{CF}}$ in Eq. (2) are defined in a “local coordinate system” convenient for the $D_{2}$ symmetry due to the oxygens surrounding the RE sites (more is discussed in Sec. III below, and further details can be found in Appendix B). Our results here make use of the parameters $\tilde{B}_{q}^{k}$ estimated via neutron spectroscopy on TGG by Wawryńczk et al. [24]. These TGG parameters are listed in the left column of Table 1 together with the corresponding values for TbIG obtained by Hau et al. via infra-red spectroscopy [20].
@@
In the mean field calculation below, the TGG parameters for $H_{\mathrm{CF}}$ in Eq. (2) are used to set the unperturbed CF Hamiltonian for Tb ions in TbIG. Each Tb site is then independently and separately examined using the Hamiltonian in Eq. (1) and one of the the six different coordinate systems exploiting Eq. (2) (details in Appendix B). We underline that our theoretical results for the SSE by CFE in TbIG are obtained by using the TGG parameters, from the left column of Table 1.
TGG, Ref. 24 TbIG, Ref. 20
@@
Table 1: Parameters (in meV) for the Stevens’ CF Hamiltonian in Eq. (2). All values are from the experimental literature: neutron spectroscopy on Tb3Ga5O12 (TGG); optical spectroscopy on Tb3Fe5O12 (TbIG) [20, 24]. For TbIG, the $\tilde{B}_{q}^{k}$ Stevens parameters (here reported in meV) are converted from the Wybourne parameters in Ref. 20 (where they are expressed in cm-1).
@@ III Mean field solution of each Tb site
where $n_{\alpha}$ and $z_{\alpha\beta}$ denote, respectively, the site number in a unit cell of the model (see Fig. 1) and the number of neighbouring sites, with $\alpha,\beta=a,d,c$ for the site species. Here, the $i$ index e.g., Eq. (6) indicates one of the six different sublattices of Tb sites, with $i=1,2,\dots,6$ following the notation adopted in Refs. 22, 23. Among these, the sites 1, 2 and 3 (4, 5, and 6) are assigned to the symmetry and g-tensors of the $\mathbf{C}$-sites ($\mathbf{C^{\prime}}$-sites) [26, 27, 29]. The three $\mathbf{C}$-sites ($\mathbf{C^{\prime}}$-sites) are related by 120 degrees rotation around the [111] axis of a garnet crystal. In Fig. 8 of the Appendix A, we show a view of the Tb atoms of a cubic unit cell in order to see this clearly, and visualize the distinction of $\mathbf{C}$ and $\mathbf{C^{\prime}}$ sites.
@@
$\cdot$ Local RE coordinates, $\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{z}_{i}$ (see Appendix B);
@@
These coordinate axes are assigned to represent features of the real garnet crystals, with the Fe moments aligning ferrimagnetically along the [111] of the “laboratory coordinate”—i.e., the azimuth $\mathbf{z}=[111]$ of the “magnetic coordinate”—and the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ symmetries dictating the “local coordinates” for the RE sublattices (see Appendix B[26, 27, 29, 22, 23, 24, 21]. We note that the definition of which rare-earth sites of the original cubic cell belong to $\mathbf{C}$ or $\mathbf{C^{\prime}}$ depends on the choice of the net-magnetic moment, here taken to be in the $[111]$ direction. If we took a domain in another equivalent direction, e.g. in the $[11{\overline{1}}]$ direction, the definitions of sites belonging to $\mathbf{C}$ or $\mathbf{C^{\prime}}$ would be permuted. In $H_{\mathrm{MFi}}$, the “magnetic coordinate” system is used even for the Tb moments, for a unified description with the Fe moments. Using Eq. (6), the CF Hamiltonian is rotated from the $i$-th “local coordinate” of the Tb sites into the “magnetic coordinate” convenient for the mean-field model of Fig.1. These rotations are realised by the Wigner matrices ${\cal D}(0,\theta,\phi),{\cal D}(\alpha_{i},\beta_{i},\gamma_{i})$, with $(0,\theta=\arctan\!\!\sqrt{2},\phi=-\pi/4)$ and $(\alpha_{i},\beta_{i},\gamma_{i})$ the Euler angles of the “magnetic” and “local” coordinates, respectively, as defined in the “laboratory coordinates” [22, 23]. The explicit parametrisation of the $\alpha_{i},\beta_{i},\gamma_{i}$ Euler angles and the respective local coordinates $\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{z}_{i}$ for each $i$-th Tb sublattice are given in Appendix B.
@@
In Fig. 2, the summations of six cases for each ion, i.e., $M_{c}^{z}=\sum_{i=1~{}6}M_{c}^{z}(i)$, $M_{d}^{z}=\sum_{i=1~{}6}M_{d}^{z}(i)$, and $M_{a}^{z}=\sum_{i=1~{}6}(2/3)M_{a}^{z}(i)$, are plotted. This corresponds to a half of a primitive cell containing six Tb sites, six $d$ sites and four $a$ sites. Here, the $x,y,z$ indexes refer to the axes of the “magnetic” coordinate system as defined above. Thus, the spins ${\vec{S}}_{a}$ and ${\vec{S}}_{d}$, as calculated in Eq. (1), and the $\vec{J}_{c}$ are expressed in this same coordinate system, as anticipated. Since the $a$ and $d$ moments of the Fe sublattices are assumed to be antiparallel along the azimuth $z$—remember this coincides with the [111] of the “laboratory coordinate system”—these will only have a net $z$ component, i.e., $M_{a}^{x}=M_{a}^{y}=M_{d}^{x}=M_{d}^{y}=0$, whereas all components of the $c$ magnetic moments (Tb sites) can be non-zero below the Néel temperature. After taking the summation, the $x$ and $y$ components of Tb sites are zero, i.e., $M_{c}^{x}=\sum_{i=1~{}6}M_{c}^{x}(i)=0$ and $M_{c}^{y}=\sum_{i=1~{}6}M_{c}^{y}(i)=0$, since three sites among the $\mathbf{C}$-sites ($\mathbf{C^{\prime}}$-sites) are related by 120 rotation around the azimuth $z$. We note that this is a simplification of the model: in the real lattice the overall magnetisation along the [111] of the crystal is stable because of the net cancellation of the transverse contribution of the c-sites, and we would also expect induced transverse components on at least some of the Fe sites, as seen in the simplest model [21].
@@
The temperature dependence of the crystal-field levels from the mean field solution of Eq. (3) are plotted in Figs. 4. The different curves can be grouped into the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ results, since the molecular field along the crystallographic [111] has the same projections on the $i=1,2$ and $3$ RE sites and, separately, on the $i=4,5$ and $6$, as discussed in Ref. 21. At the Néel temperature, the molecular field goes to zero and thus the TGG energies observed by Wawryńczk et al. are reproduced [24]. It is noteworthy that, as the molecular field increases for lower temperatures, low energy excitations appear at around 3 meV and 6 meV, as observed by Kawamoto et al. [16].
@@
Figure 5: Magnetic field (B) dependence of the lowest CFE of a $\mathbf{C}$-site at $T=$ 10 K. The horizontal axis $B$ is the applied field in the [111] direction. The inset is the same but for the second CFE. For the $\mathbf{C}$-sites the lowest energy levels are higher in energy and increase with field, (see Appendix C ).
Remarkably, the lowest CFE decreases with increasing the strength of the applied magnetic field. This is in sharp contrast to the magnetic excitation gap of a ferrimagnetic Heisenberg model [13] and illustrates the importance of including the effects of the strong anisotropies for compounds where rare-earth ions have non-zero orbital angular momentum. On the other hand, the second CFE of the $\mathbf{C}$-site (see inset of Fig. 5) increases with the magnetic field. Although the results in Fig. 5 might seem to be at odds with our intuition, the magnetization of the Fe sites resolves this apparent contradiction. The magnetization of the iron sites generates a molecular field with magnitude around 14.4 T. When an external magnetic field is applied in addition to this, the lowest CFE decreases for up to around 15 T of the external magnetic field. This can be verified already at a single-ion level: for the $\mathbf{C}$-sites the CF Hamiltonian from TGG with a magnetic field applied along the crystallographic [111] reproduces this behaviour, whilst for the same conditions for the $\mathbf{C^{\prime}}$-sites the levels all increase with magnetic field strength, closer to intuition based on isotropic models (details can be found in Appendix C). Therefore, one can say that Fig. 5 is reasonable in TbIG.
Reflecting upon these findings, a question arises: Does the low energy CFE contribute to generate a spin current? In the next section, we propose a simple model to examine this question and provide an answer. Rather than using the full set of levels, at low temperatures we can restrict our attention to the lowest one, drawn in red in Fig. 5. The quantitative feature we take from this curve is encoded as a field dependent local effective anisotropy.